OpenStudy (anonymous):

9y^5/15x^7*5x^3/45y^2

OpenStudy (anonymous):

\[\frac{ 9y ^{5} }{ 15x ^{7}}\times \frac{ 5x ^{3} }{ 45y ^{2} }\]

OpenStudy (anonymous):

http://www.wolframalpha.com/ always go to this site for math problems it is a life saver it will give you the answer and if you make an account it shows you how it got the answers step by step

OpenStudy (anonymous):

Thanks! I actually just figured it out.

OpenStudy (anonymous):

oh thats great! :) good job